# Signal attenuation enables scalable decentralized multi-agent reinforcement learning over networks

### v1

@@ Abstract
Classic multi-agent reinforcement learning (MARL) methods require that agents enjoy global state observability, preventing development of decentralized algorithms and limiting scalability. Recent work has shown that, under assumptions on decaying inter-agent influence, global observability can be replaced by local neighborhood observability at each agent, enabling decentralization and scalability. Real-world applications enjoying such decay properties remain underexplored, however, despite the fact that signal power decay, or signal attenuation, due to path loss is an intrinsic feature of many problems in wireless communications and radar networks. In this paper, we show that signal attenuation enables decentralization in MARL by considering the illustrative special case of performing power allocation for target detection in a radar network. To achieve this, we propose two new constrained multi-agent Markov decision process formulations of this power allocation problem, derive local neighborhood approximations for global value function and gradient estimates and establish corresponding error bounds, and develop decentralized saddle point policy gradient algorithms for solving the proposed problems. Our approach, though oriented towards the specific radar network problem we consider, provides a useful model for future extensions to additional problems in wireless communications and radar networks.
@@ II-A Radar network model
We consider a radar network composed of a set $\mathcal{N}=\{1,\ldots,n\}$ networked radars. In the presence of a target, the received signal for radar $i\in\mathcal{N}$ is given by [15, 16]:
$x^{i}=\alpha^{i}\sqrt{a^{i}}y^{i}+\sum_{i\in\mathcal{N}\setminus\{i\}}\beta^{ji}\sqrt{a^{j}}y^{j}+w^{i},$ (1)
where $y^{i}=\psi^{i}z^{i}$ describes the transmitted signal from radar $i$ and $z^{i}=\left[1\quad e^{j2\pi f_{D,i}}\quad\ldots\quad e^{j2\pi(N-1)f_{D,i}}\right]^{T}$ is the Doppler steering vector of radar $i$ associated with the desired target, $f_{D,i}$ denotes the normalized Doppler shift as seen by radar $i$, $N$ is the number of received pulses at each timestep, and $\psi^{i}$ denotes the predesigned waveform transmitted from radar $i$. Furthermore, the parameter $\alpha^{i}$ denotes the desired channel gain in the target direction, $a^{i}$ denotes the transmission power of radar $i$, $\beta^{ji}$ describes the cross-channel gain between radars $i$ and $j$, and $w^{i}$ denotes zero-mean white Gaussian noise.
We assume that $\alpha^{i}\sim CN(0,h^{t}_{ii})$, $\beta^{ij}\sim CN(0,c_{ij}(h^{t}_{ij}+h^{d}_{ij}))$, and $w^{i}\sim CN(0,\sigma^{2})$, where $c_{ij}h^{t}_{ij}$ represents the variance of the channel gain for the radar $i$-target-radar $j$ path, $c_{ij}h^{d}_{ij}$ represents the variance of the channel gain for the direct radar $i$-radar $j$ path, and $c_{ij}$ denotes the cross-correlation coefficient between the $i$th and $j$th radars. Note that $c_{ii}=1$, so the variance of the channel gain for the radar $i$-target-radar $i$ path is simply $h^{t}_{ii}$. The variances are given by
@@
$SINR_{i}=\frac{h^{\tau}_{ii}a^{i}}{\sigma^{2}+\sum_{j\in\mathcal{N}\setminus\{i\}}c_{ji}\left(h^{d}_{ji}+h^{\tau}_{ji}\right)a^{j}}.$ (4)
@@ II-B Constrained multi-agent MDP formulations
To capture the underlying system model described above, we use a constrained multi-agent Markov decision process (CMAMDP) $(\mathcal{S},\mathcal{A},p,\mathcal{N},\{r^{i}\}_{i\in\mathcal{N}},\{c^{i}\}_{i\in\mathcal{N}})$, defined below. Let the set of agents $\mathcal{N}=\{1,\ldots,n\}$ correspond to the $n$ radars. Let $\mathcal{S}=\mathcal{S}^{1}\times\ldots\times\mathcal{S}^{n}$ denote the joint state space, where the $i$th component $\mathcal{S}^{i}\subset\mathbb{R}^{k}\times\mathbb{R}^{m}$ corresponds to set of possible locations and movements of the target and radar $i$, i.e., $s^{i}=(s^{i}_{target},s^{i}_{radar})$ with $s^{i}_{target}\in\mathbb{R}^{k}$ and $s^{i}_{radar}\in\mathbb{R}^{m}$. When the target is moving in $\mathbb{R}^{3}$ and the radars are moving in $\mathbb{R}^{2}$, for example, we may take $k=9$ and $m=6$ and let $s^{i}_{target}$ and $s^{i}_{radar}$ correspond to the positions, velocities, and accelerations of the target and radar, respectively. Notice that, for all $i,j\in\mathcal{N}$, we have $s^{i}_{target}=s^{j}_{target}$. Given the joint state $s_{t}\in\mathcal{S}$ at time $t$, we assume that the Doppler steering vectors, predesigned waveforms, and channel gain variances corresponding to radar $i$ are provided and that the goal of radar $i$ is to determine an appropriate power level to supply given its local state information $s^{i}_{t}$. To capture this, let $\mathcal{A}=\mathcal{A}^{1}\times\ldots\times\mathcal{A}^{n}$, where $\mathcal{A}^{i}=[0,a_{max}^{i}]$ denotes the set of possible power allocations at radar $i$.
@@
Let the transition dynamics $p:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ capture the movement of the target and radars over time. We assume in this paper that the movements of the target and radars, and therefore $p$, are independent of the power allocations applied at the radars. Given joint state $s\in\mathcal{S}$ and joint power allocation $a\in\mathcal{S}$, let $r^{i}(s,a)=SINR_{i}(s,a)$ denote the SINR received at radar $i$ obtained from equation (4) by substituting the channel gains, cross-correlation coefficients, and noise corresponding to $s$ and applying allocation $a$, i.e.,
$r^{i}(s,a)=\frac{h^{\tau}_{ii}(s)a^{i}}{\sigma^{2}(s,a)+\sum_{j\in\mathcal{N}\setminus\{i\}}c_{ji}(s)\left(h^{d}_{ji}(s)+h^{\tau}_{ji}(s)\right)a^{j}}.$ (5)
@@
Finally, let the cost $c^{i}(s,a)$ denote the cost to radar $i$ of applying power level $a^{i}$. We might simply take $c^{i}(s,a)=a^{i}$, for example, but our approach accommodates more general cost structures similarly.
@@
Let a fixed communication radius $R>0$ between the radars be given, and denote the undirected communication graph between radars by $\mathcal{G}=(\mathcal{N},\mathcal{E})$, where =$\mathcal{E}=\{(i,j)\ |\ d(i,j)\leq R\}$ and $d(i,j)$ denotes the euclidean distance between radars $i$ and $j$. We assume in this paper that any movement of the radars is such that $\mathcal{G}$ is respected, i.e., that $\mathcal{E}$ remains constant. For a given positive integer $\kappa\in\mathbb{N}^{+}$, let $\mathcal{N}_{\kappa}(i)$ denote the $\kappa$-hop neighborhood of agent $i$ with respect to $\mathcal{G}$. Note that, since $\mathcal{G}$ is undirected, the $\kappa$-hop neighborhood relation is symmetric, i.e., $i\in\mathcal{N}_{\kappa}(j)$ if and only if $j\in\mathcal{N}_{\kappa}(i)$.
@@
Let $\kappa\in\mathbb{N}^{+}$ be fixed, and assume each radar $i$ has access to the state information $s^{\mathcal{N}_{\kappa}(i)}\subset\mathcal{S}^{\mathcal{N}_{\kappa}(i)}=\{s^{j}\ |\ j\in\mathcal{N}_{\kappa}(i)\}$. To each $i$, let there be associated a parameterized policy class $\{\pi^{i}_{\theta_{i}}:\mathcal{S}^{\mathcal{N}_{\kappa}(i)}\rightarrow\Delta(\mathcal{A}^{i})\}_{\theta^{i}\in\Theta^{i}}$, where $\Theta^{i}\subseteq\mathbb{R}^{d}$ is the set of permissible policy parameters, for some positive integer $d$. Denote the induced joint policy by $\pi_{\theta}(a|s)=\prod_{i\in\mathcal{N}}\pi^{i}_{\theta^{i}}(a^{i}|s^{\mathcal{N}_{\kappa}(i)})$, where $\theta=\left[(\theta^{1})^{T}\ldots(\theta^{n})^{T}\right]^{T}\in\Theta=\Theta^{1}\times\ldots\times\Theta^{n}$ is the stacked vector of each agent’s policy parameters. Given policy $\pi_{\theta}$, under this formulation
$J_{r^{i}}(\theta)=\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{T-1}r^{i}(s_{t},a_{t})\right]$ (6)
@@
captures the average expected SINR achieved at radar $i$ and
$J_{c^{i}}(\theta)=\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{T-1}c^{i}(s_{t},a_{t})\right]$ (7)
captures the average expected cost incurred at radar $i$.
@@ III Signal attenuation enables decentralization
where the local $Q$ function $Q^{f^{i}}_{\theta}(s,a)$ is defined by
@@ III-B Signal Attenuation and Decentralization
We start this section by stating several assumptions that will be need in the subsequent analysis.
@@
There exists a function $g:\mathbb{N}^{+}\times\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$, strictly increasing in both entries, such that, for all $i\in\mathcal{N}$, all $\kappa\in\mathbb{N}^{+}$, and any $j\in\mathcal{N}^{-1}_{\kappa}(i)$, we have $d_{ij}=d(i,j)\geq g(\kappa,R)$.
@@
When $\kappa=1$, taking $g(\kappa,R)=R$ is a natural choice in Assumption 1, since, for any $j\in\mathcal{N}^{-1}_{1}(i)$, we have $d_{ij}\geq g(1,R)=R$ by definition of the neighbor relation in $\mathcal{G}$ (see Section II-B). More generally, letting $g(\kappa,R)=\kappa R$ satisfies Assumption 1 for a large class of network topologies, including when the radars are arranged in a grid topology with grid cell side lengths $R$. In general, appropriate choice of $g$ excludes pathological network topologies like that of Figure 1, where, for arbitrarily large $\kappa$, there exist radars $i,j\in\mathcal{N}$ simultaneously satisfying $d_{ij}=2R$ and $j\in\mathcal{N}^{-1}_{\kappa}(i)$.
Figure 1: Example of a pathological network topology where $d_{ij}=2R$, but which can be constructed such that $j\in\mathcal{N}^{-1}_{\kappa}(i)$, for any $\kappa$. This is excluded by Assumption 1 for appropriate choices of $g$, such as $g(\kappa,R)=\kappa R$.
@@
For all $i,j\in\mathcal{N}$, we have $\inf_{s\in\mathcal{S}}R_{i}(s)\geq 1$, $\bar{\sigma}^{RCS}=\sup_{s\in\mathcal{S}}\sigma_{ij}^{RCS}(s)<\infty$ and $\bar{\sigma}^{RCS}>0$, $\bar{c}=\sup_{s\in\mathcal{S}}c_{ji}(s)<\infty$ and $\bar{c}>0$, $\underset{\bar{}}{\sigma}^{2}=\inf_{s,a}\sigma^{2}(s,a)>0$, and $\lambda,G_{t},G_{r},G_{t}^{\prime},G_{r}^{\prime}>0$.
@@
There exist $\rho\in(0,1)$ and $m\in\mathbb{R}^{+}$ such that every joint policy $\pi_{\theta}$ satisfies $d_{TV}(d_{\pi_{\theta}}^{t}(\cdot|s_{0})||d_{\theta_{t}}(\cdot))\leq m\rho^{t}$, for any $s_{0}\in\mathcal{S}$ and for all $t\geq 0$, where $d_{\pi_{\theta}}^{t}(\cdot|s_{0})$ denotes the $t$-step state occupancy measure induced by $\pi_{\theta}$ over $\mathcal{S}$ given start state $s_{0}$.
@@
There exists $\varepsilon_{\kappa}>0$ such that, for each $i\in\mathcal{N}$ and all $j\in\mathcal{N}_{\kappa}^{-1}(i)$, we have $\left\lVert\nabla_{\theta^{i}}J_{r^{i}}(\theta)\right\rVert\leq\varepsilon_{\kappa}$, for all $\theta\in\Theta$.
Assumption 1 provides a mechanism for ensuring that the radar network provides adequate coverage of the region under consideration through appropriate choice of $g$. Assumption 2 stipulates that the local costs at each radar are independent of the costs at all other radars. This is reasonable when costs depend only on local power consumption, for example. Assumption 3 provides minimal conditions ensuring that the radar range equations (2)-(3) lead to well-conditioned local rewards (5). Assumptions 4 and 5 are commonly used in the RL [19, 20, 21] and MARL [8, 9] literatures to enable analysis of policy gradient-based methods.

### v2

@@ Abstract
Multi-agent reinforcement learning (MARL) methods typically require that agents enjoy global state observability, preventing development of decentralized algorithms and limiting scalability. Recent work has shown that, under assumptions on decaying inter-agent influence, global observability can be replaced by local neighborhood observability at each agent, enabling decentralization and scalability. Real-world applications enjoying such decay properties remain underexplored, however, despite the fact that signal power decay, or signal attenuation, due to path loss is an intrinsic feature of many problems in wireless communications and radar networks. In this paper, we show that signal attenuation enables decentralization in MARL by considering the illustrative special case of performing power allocation for target detection in a radar network. To achieve this, we propose two new constrained multi-agent Markov decision process formulations of this power allocation problem, derive local neighborhood approximations for global value function and policy gradient estimates and establish corresponding error bounds, and develop decentralized saddle point policy gradient algorithms for solving the proposed problems. Our approach, though oriented towards the specific radar network problem we consider, provides a useful model for extensions to additional problems in wireless communications and radar networks.
@@ II-A Radar network model
Figure 1: Illustrations of radar network system model of Section II-A (left) and communication neighborhoods of Section II-B (right).
We consider a radar network composed of a set $\mathcal{N}=\{1,\ldots,n\}$ networked radars. In the presence of a target, the received signal for radar $i\in\mathcal{N}$ is given by [15, 16]
$x^{i}=\alpha^{i}\sqrt{a^{i}}y^{i}+\sum_{i\in\mathcal{N}\setminus\{i\}}\beta^{ji}\sqrt{a^{j}}y^{j}+\omega^{i},$ (1)
where $y^{i}=\psi^{i}z^{i}$ describes the transmitted signal from radar $i$ and $z^{i}=\left[1\quad e^{j2\pi f_{D,i}}\quad\ldots\quad e^{j2\pi(N-1)f_{D,i}}\right]^{T}$ is the Doppler steering vector of radar $i$ associated with the desired target, $f_{D,i}$ denotes the normalized Doppler shift as seen by radar $i$, $N$ is the number of received pulses at each timestep, and $\psi^{i}$ denotes the predesigned waveform transmitted from radar $i$. Furthermore, the parameter $\alpha^{i}$ denotes the desired channel gain in the target direction, $a^{i}$ denotes the transmission power of radar $i$, $\beta^{ji}$ describes the cross-channel gain between radars $i$ and $j$, and $\omega^{i}$ denotes zero-mean white Gaussian noise.
We assume that $\alpha^{i}\sim CN(0,h^{\tau}_{ii})$, $\beta^{ij}\sim CN(0,c_{ij}(h^{\tau}_{ij}+h^{d}_{ij}))$, and $\omega^{i}\sim CN(0,(\sigma^{i})^{2})$ for an $i$-dependent $\sigma^{i}>0$, where $c_{ij}h^{\tau}_{ij}$ represents the variance of the channel gain for the radar $i$-target-radar $j$ path, $c_{ij}h^{d}_{ij}$ represents the variance of the channel gain for the direct radar $i$-radar $j$ path, and $c_{ij}$ denotes the cross-correlation coefficient between the $i$th and $j$th radars. Note that $c_{ii}=1$, so the variance of the channel gain for the radar $i$-target-radar $i$ path is simply $h^{\tau}_{ii}$. The variances are given by the radar range equations [13]
@@
$SINR_{i}=\frac{h^{\tau}_{ii}a^{i}}{(\sigma^{i})^{2}+\sum_{j\in\mathcal{N}\setminus\{i\}}c_{ji}\left(h^{d}_{ji}+h^{\tau}_{ji}\right)a^{j}},$ (4)
where $\sigma^{i}$ corresponds to the noise received at radar $i$. See Figure 1 for an illustration of the radar network system model.
@@ II-B Constrained multi-agent MDP formulations
To capture the underlying system model described above, we use a constrained multi-agent Markov decision process (CMAMDP) $(\mathcal{S},\mathcal{A},p,\mathcal{N},\{r^{i}\}_{i\in\mathcal{N}},\{c^{i}\}_{i\in\mathcal{N}})$, defined below. Let the set of radars $\mathcal{N}=\{1,\ldots,n\}$ correspond to the $n$ radars. Let $\mathcal{S}=\mathcal{S}^{1}\times\ldots\times\mathcal{S}^{n}$ denote the joint state space, where the $i$th component $\mathcal{S}^{i}\subset\mathbb{R}^{k}\times\mathbb{R}^{m}$ corresponds to set of possible locations and movements of the target and radar $i$, i.e., $s^{i}=(s^{i}_{target},s^{i}_{radar})$ with $s^{i}_{target}\in\mathbb{R}^{k}$ and $s^{i}_{radar}\in\mathbb{R}^{m}$. When the target is moving in $\mathbb{R}^{3}$ and the radars are moving in $\mathbb{R}^{2}$, for example, we may take $k=9$ and $m=6$ and let $s^{i}_{target}$ and $s^{i}_{radar}$ correspond to the positions, velocities, and accelerations of the target and radar, respectively. Notice that, for all $i,j\in\mathcal{N}$, we have $s^{i}_{target}=s^{j}_{target}$. Given the joint state $s_{t}\in\mathcal{S}$ at time $t$, we assume that the Doppler steering vectors, predesigned waveforms, and channel gain variances corresponding to radar $i$ are provided and that the goal of radar $i$ is to determine an appropriate transmission power level to supply given its local state information $s^{i}_{t}$. To capture this, let $\mathcal{A}=\mathcal{A}^{1}\times\ldots\times\mathcal{A}^{n}$, where $\mathcal{A}^{i}=[0,a^{max}]$ denotes the set of possible power allocations at radar $i$, for a predetermined, finite, maximum power allocation $a^{max}>0$ over all radars.
@@
Let the transition dynamics $p:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ capture the movement of the target and radars over time. We assume in this paper that the movements of the target and radars, and therefore $p$, are independent of the transmission power allocations applied at the radars. Given joint state $s\in\mathcal{S}$ and joint power allocation $a\in\mathcal{S}$, let $r^{i}(s,a)=SINR_{i}(s,a)$ denote the SINR received at radar $i$ obtained from equation (4) by substituting the channel gains, cross-correlation coefficients, and noise corresponding to $s$ and applying allocation $a$, i.e.,
$r^{i}(s,a)=\frac{h^{\tau}_{ii}(s)a^{i}}{(\sigma^{i}(s,a))^{2}+\sum_{j\in\mathcal{N}\setminus\{i\}}c_{ji}(s)\left(h^{d}_{ji}(s)+h^{\tau}_{ji}(s)\right)a^{j}},$ (5)
where we abuse notation to let $h^{\tau}_{ii}(s),h^{\tau}_{ji}(s),h^{d}_{ji}(s),c_{ji}(s),\sigma^{i}(s,a)$, and $\sigma^{i}_{\kappa}(s,a)$ correspond to the expressions appearing in (4) when the system is in state $s$ and joint action $a$ is selected. Similarly, define the local neighborhood reward by
$r^{i}_{\kappa}(s,a)=\frac{h^{\tau}_{ii}(s)a^{i}}{(\sigma^{i}_{\kappa}(s,a))^{2}+\sum_{j\in\mathcal{N}_{\kappa}(i)\setminus\{i\}}c_{ji}(s)\left(h^{d}_{ji}(s)+h^{\tau}_{ji}(s)\right)a^{j}},$ (6)
@@
which captures the SINR received at radar $i$ originating within neighborhood $\mathcal{N}_{\kappa}(i)$, where $\sigma_{\kappa}^{i}(s,a)$ denotes the noise originating within $\mathcal{N}_{\kappa}(i)$. The expression in (6) will be crucial in the theoretical results of Section III. Finally, let the cost $c^{i}(s,a)$ denote the cost to radar $i$ of applying power level $a^{i}$. We might simply take $c^{i}(s,a)=a^{i}$, for example, but our approach accommodates general cost structures.
@@
Let a fixed communication radius $R>0$ between the radars be given, and denote the undirected communication graph between radars in state $s\in\mathcal{S}$ by $\mathcal{G}(s)=(\mathcal{N},\mathcal{E}(s))$, where $\mathcal{E}(s)=\{(i,j)\ |\ d_{ij}(s)\leq R\}$ and $d_{ij}(s)$ denotes the Euclidean distance between radars $i$ and $j$ when the system is in state $s$. We note that $R$ defines a user-specified communications neighborhood, but that it does not necessarily represent a hard limit on communications within the network. We assume in this paper that the edge set $\mathcal{E}=\mathcal{E}(s)$ remains constant, i.e., that any movement of the radars leaves the topology of the communication network unchanged, and will henceforth suppress the dependence on $s$ and simply write $\mathcal{G}=(\mathcal{N},\mathcal{E})$. For a given positive integer $\kappa\in\mathbb{N}^{+}$, let $\mathcal{N}_{\kappa}(i)$ denote the $\kappa$-hop neighborhood of radar $i$ with respect to $\mathcal{G}$. Note that, since $\mathcal{G}$ is undirected, the $\kappa$-hop neighborhood relation is symmetric, i.e., $i\in\mathcal{N}_{\kappa}(j)$ if and only if $j\in\mathcal{N}_{\kappa}(i)$. Finally, let $\mathcal{N}_{\kappa}^{-1}(i)=\mathcal{N}\setminus\mathcal{N}_{\kappa}(i)$ denote the set of all radars outside $i$’s $\kappa$-hop neighborhood. See Figure 1 for an illustration of the communication neighborhoods induced by a specified communication radius $R$ on a simple example.
@@
Let $\kappa\in\mathbb{N}^{+}$ be fixed, and assume each radar $i$ has access to the state information $s^{\mathcal{N}_{\kappa}(i)}\in\mathcal{S}^{\mathcal{N}_{\kappa}(i)}=\{s^{j}\ |\ j\in\mathcal{N}_{\kappa}(i)\}$. To each $i$, let there be associated a parameterized policy class $\{\pi^{i}_{\theta_{i}}:\mathcal{S}^{\mathcal{N}_{\kappa}(i)}\rightarrow\Delta(\mathcal{A}^{i})\}_{\theta^{i}\in\Theta^{i}}$, where $\Theta^{i}\subseteq\mathbb{R}^{d}$ is the set of permissible policy parameters, for some positive integer $d$. Denote the induced joint policy by $\pi_{\theta}(a|s)=\prod_{i\in\mathcal{N}}\pi^{i}_{\theta^{i}}(a^{i}|s^{\mathcal{N}_{\kappa}(i)})$, where $\theta=\left[(\theta^{1})^{T}\ldots(\theta^{n})^{T}\right]^{T}\in\Theta=\Theta^{1}\times\ldots\times\Theta^{n}$ is the stacked vector of each radar’s policy parameters. Given policy $\pi_{\theta}$, under this formulation,
$J_{r^{i}}(\theta)$ $=\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}_{\pi_{\theta}}\Big{[}\sum_{t=0}^{T-1}r^{i}(s_{t},a_{t})\Big{]},$ (7)
$\quad J_{c^{i}}(\theta)$ $=\lim_{T\rightarrow\infty}\frac{1}{T}\mathbb{E}_{\pi_{\theta}}\Big{[}\sum_{t=0}^{T-1}c^{i}(s_{t},a_{t})\Big{]}$ (8)
@@
capture the average expected SINR achieved and average expected cost incurred, respectively, at radar $i$.
@@ III Signal attenuation enables decentralization
where the $Q$ function $Q^{f^{i}}_{\theta}(s,a)$ corresponding to $f^{i}$ is
@@ III-B Signal Attenuation and Decentralization
We start this section by stating several assumptions that will be needed in the subsequent analysis. Assumption 1 provides a mechanism for ensuring that the radar network provides adequate coverage of the region under consideration through appropriate choice of $g$. Assumption 2 stipulates that the local costs at each radar are independent of the costs at all other radars. This is reasonable when costs depend only on local power consumption, for example. Assumption 3 provides minimal conditions ensuring that the radar range equations (2)-(3) lead to well-conditioned local rewards (5). Assumptions 4 and 5 are commonly used in the RL [19, 20, 21] and MARL [8, 9] literatures to enable analysis of policy gradient-based methods. Assumption 6 bounds the effect that changes in a given radar’s policy parameters can have on rewards received by radars outside its $\kappa$-hop neighborhood, which is reasonable when rewards decay as distance between agents increases. We suspect that Assumption 6 can be proven to hold for our setting, but leave this to future work.
@@
There exists a function $g:\mathbb{N}^{+}\times\mathbb{R}^{+}\rightarrow[1,\infty)$, strictly increasing in both entries, such that, for all $i\in\mathcal{N}$, all $\kappa\in\mathbb{N}^{+}$, and any $j\in\mathcal{N}^{-1}_{\kappa}(i)$, we have $d_{ij}(s)\geq g(\kappa,R)$, for all $s\in\mathcal{S}$.
@@
When $\kappa=1$, taking $g(\kappa,R)=R$ is a natural choice in Assumption 1, since, for any $j\in\mathcal{N}^{-1}_{1}(i)$, we have $d_{ij}(s)\geq g(1,R)=R$ by definition of the neighbor relation in $\mathcal{G}$ (see Section II-B). More generally, letting $g(\kappa,R)=\kappa R$ satisfies Assumption 1 for a large class of network topologies, such as when the radars are static and arranged in a grid topology with grid cell side lengths $R$.
@@
For all $i,j\in\mathcal{N}$, we have $\inf_{s\in\mathcal{S}}R_{i}(s)\geq 1$, $\bar{\sigma}^{RCS}=\sup_{s\in\mathcal{S}}\sigma_{ij}^{RCS}(s)<\infty$ and $\bar{\sigma}^{RCS}>0$, $\bar{c}=\sup_{s\in\mathcal{S}}c_{ji}(s)<\infty$ and $\bar{c}>0$, $\underset{\bar{}}{\sigma}=\min\{\inf_{s,a}\sigma^{i}(s,a),\inf_{s,a}\sigma_{\kappa}^{i}(s,a)\}>0$, and $\lambda,G_{t},G_{r},G_{t}^{\prime},G_{r}^{\prime}>0$.
@@
There exist $\rho\in(0,1)$ and $m\in\mathbb{R}^{+}$ such that every joint policy $\pi_{\theta}$ satisfies $d_{TV}(d_{\pi_{\theta}}^{t}(\cdot|s_{0})||d_{\theta}(\cdot))\leq m\rho^{t}$, for any $s_{0}\in\mathcal{S}$ and for all $t\geq 0$, where $d_{TV}(q(\cdot)||q^{\prime}(\cdot))=\sup_{A}\lvert\int_{A}q(x)\ dx-\int_{A}q^{\prime}(x)\ dx\rvert$ denotes the total variation distance between densities $q,q^{\prime}$, and $d_{\pi_{\theta}}^{t}(\cdot|s_{0})$ denotes the $t$-step state occupancy measure induced by $\pi_{\theta}$ over $\mathcal{S}$ given start state $s_{0}$.
@@
There exists $\varepsilon_{\kappa}>0$ such that, for each $i\in\mathcal{N}$ and all $j\in\mathcal{N}_{\kappa}^{-1}(i)$, we have $\left\lVert\nabla_{\theta^{i}}J_{r^{j}}(\theta)\right\rVert\leq\varepsilon_{\kappa}$, for all $\theta\in\Theta$.